# Federated Deep Reinforcement Learning for Privacy-Preserving Robotic-Assisted Surgery

### v1

@@ Abstract
The integration of Reinforcement Learning (RL) into robotic-assisted surgery (RAS) holds significant promise for advancing surgical precision, adaptability, and autonomous decision-making. However, the development of robust RL models in clinical settings is hindered by key challenges, including stringent patient data privacy regulations, limited access to diverse surgical datasets, and high procedural variability. To address these limitations, this paper presents a Federated Deep Reinforcement Learning (FDRL) framework that enables decentralized training of RL models across multiple healthcare institutions without exposing sensitive patient information. A central innovation of the proposed framework is its dynamic policy adaptation mechanism, which allows surgical robots to select and tailor patient-specific policies in real-time, thereby ensuring personalized and Optimised interventions. To uphold rigorous privacy standards while facilitating collaborative learning, the FDRL framework incorporates secure aggregation, differential privacy, and homomorphic encryption techniques. Experimental results demonstrate a 60% reduction in privacy leakage compared to conventional methods, with surgical precision maintained within a 1.5% margin of a centralized baseline. This work establishes a foundational approach for adaptive, secure, and patient-centric AI-driven surgical robotics, offering a pathway toward clinical translation and scalable deployment across diverse healthcare environments.
Index Terms:
Federated Deep Reinforcement Learning, Autonomous Surgical Robots, Task-Based Privacy Preservation, Federated Learning, Differential Privacy, Secure Reinforcement Learning, Homomorphic Encryption, and Secure Aggregation.
@@ I-A Background and Motivation
Robotic-assisted surgery has revolutionized modern medicine, offering a paradigm shift from traditional open surgery to minimally invasive procedures. This transition has led to significant advancements, including enhanced surgical precision, diminished patient trauma, reduced post-operative complications, and accelerated recovery times [1, 2, 3]. Augmenting Robotic-Assisted Surgery (RAS) platforms with Artificial Intelligence (AI) is the next frontier, promising to further enhance surgical capabilities and autonomy. Specifically, the integration of AI can enable surgical robots to perform complex tasks with greater dexterity, adapt to unforeseen intraoperative events, and provide surgeons with real-time decision support [4, 5, 6, 7].
Within the pantheon of AI methodologies, Reinforcement Learning (RL) has emerged as a particularly potent approach for endowing surgical robots with intelligent decision-making capabilities. RL empowers autonomous agents, in this case, surgical robots, to learn optimal sequences of actions through iterative interaction with a dynamic environment, guided by reward signals [5]. RL algorithms, when utilizing real-time intraoperative information alongside historical procedural data, can empower surgical robots to enhance their control strategies, customize interventions to suit each patient’s unique anatomical and physiological characteristics, and adjust in real-time to the unpredictable and variable nature of surgical procedures [6, 7]. This capability is crucial for navigating the nuanced and often complex landscape of surgical interventions.
@@ I-B Challenges in RL-based Surgical Robotics
Despite the transformative potential of RL in RAS, several formidable challenges impede its widespread adoption and clinical translation. A primary obstacle is the inherent heterogeneity of surgical environments. These environments are characterized by significant inter-patient anatomical variability, diverse patient comorbidities, and surgeon-specific procedural preferences, all of which contribute to a high degree of complexity and pose a substantial challenge to the generalizability of RL models [8, 9, 10]. Furthermore, the development of robust RL models necessitates access to large, diverse datasets of surgical procedures. However, individual healthcare institutions often suffer from data scarcity, which limits the ability of RL models trained in isolation to generalize effectively to real-world clinical settings, particularly when encountering rare pathologies or unanticipated procedural complexities [11].
@@
Traditionally, RL-driven surgical systems have predominantly relied on centralized training paradigms, wherein sensitive patient data from multiple institutions are aggregated and stored in a central repository for model development [14]. This centralized approach introduces significant privacy risks, increasing vulnerabilities to data breaches, unauthorized access, and sophisticated attacks such as model inversion and membership inference attacks, which can expose sensitive patient information [15]. Additionally, centralized models may fail to adequately capture the nuances of institution-specific surgical practices and procedural variations, thus limiting their translational efficacy and hindering personalized surgical decision-making [16].
@@ I-C Federated Learning for Privacy-Preserving Collaboration
Federated Learning (FL) has emerged as a promising distributed learning paradigm that addresses the privacy challenges associated with centralized RL training. FL enables collaborative model training across multiple geographically distributed hospitals or healthcare institutions without the need for direct sharing of sensitive patient data [17, 18, 19]. In FL, each participating institution trains AI models locally on its private dataset. Subsequently, instead of sharing raw data, institutions share only aggregated model updates, such as gradients or parameter differentials, with a central server or aggregator. This process preserves data privacy and security by ensuring that sensitive patient information remains within the confines of individual institutions. Consequently, FL promotes robust collaborative learning across diverse clinical environments while mitigating privacy risks.
@@ I-D Proposed Federated Deep Reinforcement Learning (FDRL) Framework
To address these pressing challenges data scarcity, privacy concerns, and procedural variability we propose a novel Federated Deep Reinforcement Learning (FDRL) framework designed to enhance both security and adaptability in RAS. Motivated by these privacy imperatives and the need for enhanced adaptability and robustness in surgical RL, this research introduces a novel FDRL framework explicitly designed for RAS. Our approach integrates advanced cryptographic privacy-enhancing technologies (PETs), including differential privacy, Secure Aggregation, and Homomorphic Encryption (HE), to provide robust guarantees of patient data confidentiality throughout the federated training process [20]. A key innovation of our framework is the dynamic policy adaptation mechanism. This mechanism empowers surgical robots to intelligently select and execute the most appropriate RL policy in real-time, based on the dynamic and evolving context of the surgical procedure, including patient-specific conditions, surgical complexity, and procedural demands [21]. Through this dynamic adaptation, the proposed framework significantly enhances surgical adaptability, precision, and patient safety by leveraging a diverse repertoire of federated-trained RL policies.
@@ I-F Paper Organisation
The remainder of this paper is structured as follows. Section II provides a detailed exposition of the proposed FDRL framework’s architecture, thoroughly explaining the mechanisms for dynamic policy selection and privacy preservation. Section III presents a comparative privacy analysis between federated and centralized RL frameworks, including elaboration on privacy metrics. Section IV outlines implementation details and methodologies related to HE and Secure Aggregation. Finally, Section V concludes the paper and outlines potential avenues for future research.
@@ II Federated Reinforcement Learning Framework for Surgical Robots
Each hospital trains RL policies tailored to specific surgical procedures, such as colonoscopy or minimally invasive cardiac surgery. Given that multiple policies may exist for the same surgical task across different hospitals, a selection mechanism is required. The proposed framework evaluates available policies based on cumulative reward and predefined surgical performance metrics, ensuring that the policy demonstrating superior performance is selected for real-time execution. This dynamic selection process Optimises surgical precision and adaptability. The proposed RL framework is formulated as a Markov Decision Processes (MDP), defined by the tuple $(S,A,P,R,\gamma)$, where $S$ represents the state space, encompassing patient-specific parameters, surgical conditions, and real-time sensor inputs. $A$ denotes the action space, consisting of robotic movements, tool manipulations, and incision strategies. $P(s^{\prime}|s,a)$ is the transition probability function governing state transitions based on the applied action.
The RL objective is to determine an optimal policy $\pi^{*}(a\mid s)$ that maximizes the expected cumulative reward
$J(\pi)=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t})\right]$ (1)
@@
The robot’s actions include tool movements, incisions, suturing, etc. The robot does not know the optimal actions initially but gradually learns by experimenting with different actions, observing their outcomes (e.g., successful incision, minimal damage to tissue, or better healing outcomes), and adjusting its strategy based on these results. Each hospital or centre can train its RL model using local patient data (e.g., from its surgeries) and share model updates (e.g., gradients or weights) with a central server. The server aggregates these updates into a global model, which is then sent back to the hospitals for further improvement. The key advantage here is that the data never leaves the local institution, ensuring privacy and security, but the model is still able to learn from a large, diverse set of data across multiple hospitals.
A detailed summary of the parameters and notation used throughout this study is provided in Table I, ensuring clarity and consistency in defining key mathematical formulations and system components.
TABLE I: Comprehensive Table of Parameters and Notations
Symbol Definition
$S$ State space in MDP
$A$ Action space in MDP
$P(s^{\prime}|s,a)$ Transition probability function
$R(s,a)$ Reward function
$\gamma$ Discount factor
$\pi^{*}(a|s)$ Optimal policy
$J(\pi)$ Expected cumulative reward
$\theta_{i}$ Local model parameters at hospital $i$
$\theta$ Global model parameters
$n_{i}$ Number of training samples at hospital $i$
$n$ Total number of training samples
$L(\theta_{i};D_{i})$ Local loss function
$\eta$ Learning rate
$\lambda$ Regularization term
$\alpha_{adapt}$ Policy adaptation rate
$\Delta s_{t}$ State variation at time $t$
$\theta_{th}$ Threshold for state variation
$\sigma^{2}$ Variance of Gaussian noise (for DP)
$Enc()$ Encryption function
$A_{task}$ Task-specific accuracy
$R_{mit}$ Surgical risk mitigation score
$D_{t}$ Combined risk factor at time $t$
$F_{t}$ Force applied at time $t$
$T_{d,t}$ Tissue damage at time $t$
$C_{t}$ Critical error indicator at time $t$
$w_{1},w_{2},w_{3}$ Weighting coefficients
$d(s_{t},s_{t-1})$ Change in patient state
$\lambda_{1},\lambda_{2},\lambda_{3}$ Weighting coefficients for metrics
$I(W;D)$ MI between weights and data
$H(D)$ Entropy of dataset
$D_{KL}$ KL divergence
$\epsilon$ Privacy budget in DP
@@
One of the exciting opportunities that FL offers in this domain is the ability to train multiple policies (i.e., different RL models) that specialize in different surgical contexts or conditions. For example, one policy might be particularly good for handling minimally invasive surgery while another might be better suited for open surgery or tissue repair. Another policy could specialize in robotic-assisted precision surgeries. The global model, which combines knowledge from all hospitals, can then intelligently select and apply the appropriate policy depending on the context of the current surgery (e.g., the type of procedure being performed, the patient’s condition, or the tools available).
@@
FL ensures decentralized training while preserving privacy. We use the Federated Averaging (FedAvg) algorithm, where each hospital $i$ updates its local model $\theta_{i}$ and transmits gradients to the global model $\theta$
$\theta\leftarrow\sum_{i=1}^{N}\frac{n_{i}}{n}\theta_{i}$ (2)
@@
where $n_{i}$ is the number of training samples at hospital $i$ and $n$ is the total data across all institutions. To mitigate non-IID data challenges, we introduce weighted local updates
$\theta_{i}\leftarrow\theta_{i}-\eta\nabla L(\theta_{i};D_{i})+\lambda(\theta-\theta_{i})$ (3)
@@
where $L(\theta_{i};D_{i})$ is the local loss function, $\eta$ is the learning rate, and $\lambda$ is the regularization term.
@@ II-A Evaluation Metrics
The performance of the FDRL framework is evaluated based on Surgical Precision, which is measured via incision accuracy, tool placement, and minimal tissue damage, as well as Training Efficiency, which assesses convergence time and computational resource utilization. Adaptability is evaluated through the policy switching rate in response to dynamic patient conditions, expressed as
$\alpha_{adapt}=\frac{\sum_{t=1}^{T}\mathbb{I}(\Delta s_{t}>\theta)}{T}$ (4)
@@
$\theta^{\prime}_{i}=\theta_{i}+\mathcal{N}(0,\sigma^{2})$ (5)
@@
$\sum_{i=1}^{N}Enc(\theta_{i})=Enc\left(\sum_{i=1}^{N}\theta_{i}\right)$ (6)
preventing unauthorized access to local model updates. To facilitate real-time adaptation, we introduce a policy selection mechanism based on a meta-learning approach
$\pi^{*}(s)=\arg\max_{\pi_{i}}\mathbb{E}[J(\pi_{i})|s]$ (7)
@@
Figure 2: Federated Deep Reinforcement Learning Architecture for Privacy-Preserving Robotic-Assisted Surgery. The framework illustrates a multi-layered workflow comprising local RL training at distributed hospitals usingDP and HE, secure aggregation at the federated aggregator, and optimal surgical policy selection at the robotic decision layer using a MSS.
Fig. 2 depicts the FDRL workflow, where hospitals train local RL policies on private data with DP noise and HE encryption. Encrypted updates are securely aggregated by the FL aggregator and distributed as a global model. A surgical robot evaluates policies using MSS, enabling adaptive selection of the optimal RL policy for precision and privacy in RAS.

### v2

@@ Abstract
The integration of Reinforcement Learning (RL) into robotic-assisted surgery (RAS) holds significant promise for advancing surgical precision, adaptability, and autonomous decision-making. However, the development of robust RL models in clinical settings is hindered by key challenges, including stringent patient data privacy regulations, limited access to diverse surgical datasets, and high procedural variability. To address these limitations, this paper presents a Federated Deep Reinforcement Learning (FDRL) framework that enables decentralised training of RL models across multiple healthcare institutions without exposing sensitive patient information. A central innovation of the proposed framework is its dynamic policy adaptation mechanism, which allows surgical robots to select and tailor patient-specific policies in real-time, thereby ensuring personalised and optimised interventions. To uphold rigorous privacy standards while facilitating collaborative learning, the FDRL framework incorporates secure aggregation, differential privacy, and homomorphic encryption techniques. Experimental results demonstrate a 60% reduction in privacy leakage compared to conventional methods, with surgical precision maintained within a 1.5% margin of a centralised baseline. This work establishes a foundational approach for adaptive, secure, and patient-centric AI-driven surgical robotics, offering a pathway toward clinical translation and scalable deployment across diverse healthcare environments.
@@ I-A Background and Motivation
Robotic-assisted surgery has revolutionised modern medicine, offering a paradigm shift from traditional open surgery to minimally invasive procedures. This transition has led to significant advancements, including enhanced surgical precision, diminished patient trauma, reduced post-operative complications, and accelerated recovery times [1, 2, 3]. Augmenting Robotic-Assisted Surgery (RAS) platforms with Artificial Intelligence (AI) is the next frontier, promising to further enhance surgical capabilities and autonomy. Specifically, the integration of AI can enable surgical robots to perform complex tasks with greater dexterity, adapt to unforeseen intraoperative events, and provide surgeons with real-time decision support [4, 5, 6, 7].
Within the pantheon of AI methodologies, Reinforcement Learning (RL) has emerged as a particularly potent approach for endowing surgical robots with intelligent decision-making capabilities. RL empowers autonomous agents, in this case, surgical robots, to learn optimal sequences of actions through iterative interaction with a dynamic environment, guided by reward signals [5]. RL algorithms, when utilising real-time intraoperative information alongside historical procedural data, can empower surgical robots to enhance their control strategies, customise interventions to suit each patient’s unique anatomical and physiological characteristics, and adjust in real-time to the unpredictable and variable nature of surgical procedures [6, 7]. This capability is crucial for navigating the nuanced and often complex landscape of surgical interventions.
@@ I-B Challenges in RL-based Surgical Robotics
Despite the transformative potential of RL in RAS, several formidable challenges impede its widespread adoption and clinical translation. A primary obstacle is the inherent heterogeneity of surgical environments. These environments are characterised by significant inter-patient anatomical variability, diverse patient comorbidities, and surgeon-specific procedural preferences, all of which contribute to a high degree of complexity and pose a substantial challenge to the generalisability of RL models [8, 9, 10]. Furthermore, the development of robust RL models necessitates access to large, diverse datasets of surgical procedures. However, individual healthcare institutions often suffer from data scarcity, which limits the ability of RL models trained in isolation to generalise effectively to real-world clinical settings, particularly when encountering rare pathologies or unanticipated procedural complexities [11].
@@
Traditionally, RL-driven surgical systems have predominantly relied on centralised training paradigms, where sensitive patient data from multiple institutions are aggregated and stored in a central repository for model development [14]. This centralised approach introduces significant privacy risks, increasing vulnerabilities to data breaches, unauthorised access, and sophisticated attacks such as model inversion and membership inference attacks, which can expose sensitive patient information [15]. Additionally, centralised models may fail to adequately capture the nuances of institution-specific surgical practices and procedural variations, thus limiting their translational efficacy and hindering personalised surgical decision-making [16].
@@ I-C Federated Learning for Privacy-Preserving Collaboration
Federated Learning (FL) has emerged as a promising distributed learning paradigm that addresses the privacy challenges associated with centralised RL training. FL enables collaborative model training across multiple geographically distributed hospitals or healthcare institutions without the need for direct sharing of sensitive patient data [17, 18, 19]. In FL, each participating institution trains AI models locally on its private dataset. Subsequently, instead of sharing raw data, institutions share only aggregated model updates, such as gradients or parameter differentials, with a central server or aggregator. This process preserves data privacy and security by ensuring that sensitive patient information remains within the confines of individual institutions. Consequently, FL promotes robust collaborative learning across diverse clinical environments while mitigating privacy risks.
@@ I-D Proposed Federated Deep Reinforcement Learning (FDRL) Framework
To address these pressing challenges, including data scarcity, privacy concerns, and procedural variability we propose a novel Federated Deep Reinforcement Learning (FDRL) framework designed to enhance both security and adaptability in RAS. Motivated by these privacy imperatives and the need for enhanced adaptability and robustness in surgical RL, this research introduces a novel FDRL framework explicitly designed for RAS. Our approach integrates advanced cryptographic privacy-enhancing technologies (PETs), including differential privacy, Secure Aggregation, and Homomorphic Encryption (HE), to provide robust guarantees of patient data confidentiality throughout the federated training process [20]. A key innovation of our framework is the dynamic policy adaptation mechanism. This mechanism empowers surgical robots to intelligently select and execute the most appropriate RL policy in real-time, based on the dynamic and evolving context of the surgical procedure, including patient-specific conditions, surgical complexity, and procedural demands [21]. Through this dynamic adaptation, the proposed framework significantly enhances surgical adaptability, precision, and patient safety by leveraging a diverse repertoire of federated-trained RL policies.
@@ I-F Paper Organisation
The remainder of this paper is structured as follows. Section II provides a detailed exposition of the proposed FDRL framework’s architecture, thoroughly explaining the mechanisms for dynamic policy selection and privacy-preservation. Section III presents a comparative privacy analysis between federated and centralised RL frameworks, including elaboration on privacy metrics. Section IV outlines implementation details and methodologies related to HE and Secure Aggregation. Finally, Section V concludes the paper and outlines potential avenues for future research.
Ethical Statement
All datasets used in this study were synthetically generated for research purposes; no real patient data were employed.
@@ II Federated Reinforcement Learning Framework for Surgical Robots
Each hospital trains RL policies tailored to specific surgical procedures, such as colonoscopy or minimally invasive cardiac surgery. Given that multiple policies may exist for the same surgical task across different hospitals, a selection mechanism is required. The proposed framework evaluates available policies based on cumulative reward and predefined surgical performance metrics, ensuring that the policy demonstrating superior performance is selected for real-time execution. This dynamic selection process optimises surgical precision and adaptability. The proposed RL framework is formulated as a Markov Decision Process (MDP), defined by the tuple $(S,A,P,R,\gamma)$, where $S$ represents the state space, encompassing patient-specific parameters, surgical conditions, and real-time sensor inputs. $A$ denotes the action space, consisting of robotic movements, tool manipulations, and incision strategies. $P(s^{\prime}|s,a)$ is the transition probability function governing state transitions based on the applied action.
The RL objective is to determine an optimal policy $\pi^{*}(a\mid s)$ that maximises the expected cumulative reward
$J(\pi)=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t})\right].$ (1)
@@
The robot’s actions include tool movements, incisions, suturing, etc. Initially unaware of optimal actions, the robot gradually learns through experience with different actions, observing their outcomes (e.g., successful incision, minimal damage to tissue, or better healing outcomes), and adjusting its strategy based on these results. Each hospital or centre can train its RL model using local patient data (e.g., from its surgeries) and share model updates (e.g., gradients or weights) with a central server. The server aggregates these updates into a global model, which is then sent back to the hospitals for further improvement. The key advantage here is that the data never leaves the local institution, ensuring privacy and security, but the model is still able to learn from a large, diverse set of data across multiple hospitals. As summarised in Table I, the key parameters include local and global policies, privacy metrics, and surgical performance indicators.
TABLE I: Grouped Parameters and Notations Used in the Proposed Framework
Symbol Definition [Scope, Unit/Type]
Reinforcement Learning and MDP Terms
$S$ State space in MDP [global, categorical]
AA Action space in MDP [global, categorical]
$P(s^{\prime}|s,a)$ State transition probability [global, probability]
$R(s,a)$ Reward function [global, scalar]
$\gamma$ Discount factor [global, unitless]
$\pi^{*}(a|s)$ Optimal policy [global, probability distribution]
$J(\pi)$ Expected cumulative reward [global, scalar]
$\boldsymbol{\theta}_{i}$ Local model parameters at hospital $i$ [local, vector]
$\boldsymbol{\theta}$ Global model parameters [global, vector]
$L(\boldsymbol{\theta}_{i};D_{i})$ Local loss function [local, scalar]
$\eta$ Learning rate [global, unitless]
$\lambda$ Regularisation coefficient [global, unitless]
$\alpha_{adapt}$ Policy adaptation rate [global, ratio]
$\Delta s_{t}$ State variation at time $t$ [local, variable-specific]
$\theta_{th}$ State variation threshold [global, variable-specific]
Privacy and Security Terms
$\epsilon$ Privacy budget in DP [global, unitless]
$\sigma^{2}$ Variance of Gaussian noise [global, variance]
$Enc()$ Encryption function [global, functional]
$I(W;D)$ Mutual information between weights and data [global, bits]
$H(D)$ Entropy of dataset [global, bits]
$D_{KL}$ KL divergence [local, unitless]
Surgical Contextual Metrics
$A_{task}$ Task-specific accuracy [local, ratio]
$R_{mit}$ Surgical risk mitigation score [local, score]
$D_{t}$ Combined risk factor at time $t$ [local, score]
$F_{t}$ Force applied at time $t$ [local, Newtons (N)]
$T_{d,t}$ Tissue damage at time $t$ [local, damage index]
$C_{t}$ Critical error indicator at time $t$ [local, binary (0/1)]
$d(s_{t},s_{t-1})$ Change in patient state [local, variable-specific]
Weighting Coefficients and General Terms
$n_{i}$ Number of training samples at hospital $i$ [local, integer]
$n$ Total number of training samples [global, integer]
$w_{1},w_{2},w_{3}$ Metric weighting coefficients [global, unitless]
$\lambda_{1},\lambda_{2},\lambda_{3}$ Metric-specific weights [global, unitless]
@@
One of the exciting opportunities that FL offers in this domain is the ability to train multiple policies (i.e., different RL models) that specialize in different surgical contexts or conditions. For example, one policy might be particularly good for handling minimally invasive surgery while another might be better suited for open surgery or tissue repair. Another policy could specialise in robotic-assisted precision surgeries. The global model, which combines knowledge from all hospitals, can then intelligently select and apply the appropriate policy depending on the context of the current surgery (e.g., the type of procedure being performed, the patient’s condition, or the tools available).
@@
FL ensures decentralised training while preserving privacy. We use the Federated Averaging (FedAvg) algorithm, where each hospital $i$ updates its local model $\theta_{i}$ and transmits gradients to the global model $\theta$.
$\theta\leftarrow\sum_{i=1}^{N}\frac{n_{i}}{n}\theta_{i}.$ (2)
@@
where $n_{i}$ is the number of training samples at hospital $i$ and $n$ is the total data across all institutions. To mitigate non-IID data challenges, we introduce weighted local updates.
$\theta_{i}\leftarrow\theta_{i}-\eta\nabla L(\theta_{i};D_{i})+\lambda(\theta-\theta_{i}).$ (3)
@@
where $L(\theta_{i};D_{i})$ is the local loss function, $\eta$ is the learning rate, and $\lambda$ is the regularisation term.
@@ II-A Evaluation Metrics
The performance of the FDRL framework is evaluated based on Surgical Precision, which is measured via incision accuracy, tool placement, and minimal tissue damage, as well as Training Efficiency, which assesses convergence time and computational resource utilisation. Adaptability is evaluated through the policy switching rate in response to dynamic patient conditions, expressed as
$\alpha_{adapt}=\frac{\sum_{t=1}^{T}\mathbb{I}(\Delta s_{t}>\theta)}{T}.$ (4)
@@
$\theta^{\prime}_{i}=\theta_{i}+\mathcal{N}(0,\sigma^{2}).$ (5)
@@
$\sum_{i=1}^{N}Enc(\theta_{i})=Enc\left(\sum_{i=1}^{N}\theta_{i}\right).$ (6)
preventing unauthorized access to local model updates. To facilitate real-time adaptation, we introduce a policy selection mechanism based on a meta-learning approach.
$\pi^{*}(s)=\arg\max_{\pi_{i}}\mathbb{E}[J(\pi_{i})|s].$ (7)
@@
Figure 2: Federated Deep Reinforcement Learning architecture for privacy-preserving robotic-assisted surgery. The framework consists of decentralised RL training at hospital nodes using differential privacy (DP) and homomorphic encryption (HE), secure aggregation at the federated server, and adaptive policy selection at the robotic decision-making layer.
Fig. 2 depicts the FDRL workflow, where hospitals train local RL policies on private data with DP noise and HE encryption. Encrypted updates are securely aggregated by the FL aggregator and distributed as a global model. A surgical robot evaluates policies using Multi-Stage Selection (MSS), enabling adaptive selection of the optimal RL policy for precision and privacy in RAS.